### v2

# Evaluation of machine-learning models to measure individualized treatment effects from randomized clinical trial data with time-to-event outcomes

Abstract
Objective: In randomized clinical trials, prediction models can be used to explore the relationships between patients’ variables (e.g., clinical, pathological, or lifestyle variables, and also biomarker or genomic data) and treatment effect magnitude. Our aim was to evaluate flexible machine learning models capable of incorporating interactions and nonlinear effects from high-dimensional data to estimate individualized treatment recommendations in trials with time-to-event outcomes.
Methods: We compared survival models based on neural networks (CoxCC and CoxTime) and random survival forests (Interaction Forests) against a Cox proportional hazards model with an adaptive LASSO (ALASSO) penalty as a benchmark. For individualized treatment recommendations in the survival setting, we adapted metrics originally designed for binary outcomes to accommodate time-to-event data with censoring. These adapted metrics included the C-for-Benefit, the E50-for-Benefit, and the root mean squared error for treatment benefit. An extensive simulation study was conducted using two different data generation processes incorporating nonlinearity and interactions. The models were applied to gene expression and clinical data from three cancer clinical trial data sets.
Results: In the first data generation process, neural networks outperformed ALASSO in terms of calibration while the Interaction Forests showed superior C-for-benefit performance. In the second data generation process, both machine learning methods outperformed the benchmark linear ALASSO method across discrimination, calibration, and RMSE metrics. In the cancer trial data sets, the machine learning methods often performed better than ALASSO, particularly IF in terms of C-for-benefit, and either a neural network or IF for calibration measures addressing treatment benefit. Conclusion: Machine learning-based survival models can efficiently estimate conditional average treatment effects from randomized trials when strong nonlinear and interaction effects are expected. These flexible approaches offer valuable alternatives to traditional regression methods for developing treatment recommendation rules from clinical trials with time-to-event outcomes and medium to high-dimensional data set sizes.
Keywords Randomized Control Trial, Feedforward Neural Network, Interaction Forest, LASSO, Personalized Medicine, Individualized Treatment Effect
1 Introduction
Precision medicine involves identifying the intervention that is most likely to be beneficial for a given patient. It is defined by the European Society for Medical Oncology (Yates et al., 2018) as a "healthcare approach with the primary aim of identifying which interventions are likely to be of most benefit for which patients based upon the features of the individual and their disease". It focuses on characterizing the variability in patients’ response to treatment and estimating treatment benefits.
In randomized controlled trials (RCTs), in which patients are randomly assigned to two treatment groups: one that receives the experimental treatment, often called the treatment group, and a second group that does not receive the treatment, often called the control group, the treatment benefit is typically calculated as an average across the study population. In an RCT, an individual treatment benefit can be defined as the difference in survival outcomes for a given patient at a pre-specified time point with and without the treatment. The assessment of this treatment benefit is based on a counterfactual approach (Rubin, 1974), as both outcomes cannot be observed for a given individual.
In this work, we explore the potential of two families of machine learning methods to estimate conditional average treatment effects from data in a medium to high-dimensional setting and in the context of an RCT, aiming to predict invidualized treatment effects. We apply $2$ strategies for the feedforward neural networks (FNNs) based on a specific loss function in a continuous time framework (CoxCC and CoxTime) (Kvamme et al., 2019), and Interaction Forests (IF), an algorithm based on diversity forests that accounts for bidirectional interactions. We benchmark these methods with a penalized Cox Proportional Hazards (CoxPH) model (Cox, 1972) with an adaptive LASSO penalty, as previously recommended by Ternès et al. (2017) for this setting.
Our study examines the potential of these methods to estimate individual patient treatment benefits from data in a high-dimensional setting, where nonlinear and interaction effects are present. We apply calibration and discrimination measures specifically designed for the potential outcome framework and adapt some of these measures to a time-to-event outcome.
The performances of the different models are evaluated and compared using a simulation study with two different data generation processes, which simulate data with the presence of nonlinearity and interactions. We also present the results of two case studies from clinical trials in cancer utilizing gene expression data.
2 Related work
In an RCT, the main analysis focuses on the average treatment effect, which refers to the impact of a treatment on a specific outcome variable. This outcome can be of various types, such as a binary outcome or time-to-event. In this work, we are interested in the heterogeneity of the treatment effect based on patient characteristics. An overview of recent developments in statistical methods for assessing the heterogeneity of treatment effects, with a particular focus on subgroup identification and the estimation of individualized treatment strategies is provided in Lipkovich et al. (2024). For survival outcomes, the FNN model DeepSurv, Katzman et al. (2018) provide a personalized recommendation method derived from randomized data, which may predict survival in a subgroup of patients. Klaveren et al. (2017) focus on outcome risk and define treatment benefit as the difference between outcome risk with and without therapy. They define treatment benefit in $2$ different ways. Observed treatment benefit is the difference in survival outcome between $2$ patients with the same predicted benefit but with different treatment assignments. The predicted treatment benefit is the predicted risk with a treatment for a given patient minus the predicted risk with an alternative treatment option for the same patient.
In the counterfactual framework, a matching procedure should be implemented to compute the observed treatment benefit. Different methods can create pairs of patients with similar profiles but different treatment assignments. Patients can be matched on predicted benefit (Klaveren et al., 2017): $2$ patients are paired if they have close predicted benefit and discordant treatment assignments. Klaveren et al. (2017) compare this procedure with an alternative matching based on covariates. This second matching procedure is also applied by Maas et al. (2023). They use the Mahalanobis distance between patients’ characteristics to create pairs of patients.
Klaveren et al. (2017) study binary outcomes and predicted risk, discussing how to extend their measure to time-to-event data. Rekkas et al. (2023) estimate individualized treatment effects from an RCT, comparing various risk-based approaches with a binary outcome. Bouvier et al. (2024) investigates methods for evaluating individualized treatment effects with time-to-event outcomes using individual participant data meta-analysis, highlighting the potential of multiple clinical trial data sets in a meta-analysis to predict individualized treatment effects.
3 Material and Methods
3.1 Notations
Let the random variable $T\in\mathbb{R}^{+}$ represent the survival time, that is, the time between the starting point and the occurrence of a given event (e.g., the time between a patient’s treatment assignment and death). The survival function at time $t$, $S(t)$, is defined as:
$S(t)=P(T>t).$
The survival function can be obtained through the cumulative hazard $H(t)$:
$S(t)=\exp(-H(t)),$
with $H(t)=\int_{0}^{t}h(s)ds$ , and $h(t)$ the hazard rate. $H_{0}$ is the cumulative hazard at baseline for time $t=0$.
Often, $T$ is not observed for all individuals: time-to-event data is censored. Survival data are said to be right-censored at the last date of follow-up, meaning that no event has occurred. In the case of right-censoring, we define $(C_{i})_{i=1,\dots,n}$ the independent and identically distributed censoring times of $n$ individuals. We can set $\{\tilde{T}_{i},X_{i},D_{i}\}_{i=1,\dots,n}$ where $\tilde{T}=\min(T_{i},C_{i})$ is the time until death or censoring, $X_{i}=(X_{i,1},\dots,X_{i,p})^{T}\in\mathbb{R}^{p}$ denotes a $p$-dimensional vector of variables, and $D_{i}=\mathbb{I}\{\tilde{T}_{i}=T_{i}\}$ is the censoring indicator. We suppose that the survival time $T_{i}$ is independent of the censoring time conditionally on the vector of variables $X_{i}$ for $i=1,\dots,n$. $G(t)$ can be defined as the estimate of the censoring survival function:
$G(t)=P(C>t).$
3.2 Models
In this paper, we compare $3$ survival-adapted machine learning algorithms and investigate their ability to estimate conditional average treatment effects. The treatment effect can be defined as the effect of treatment $\tau$ on the outcome variable $Y$. Since $Y$ is the observed outcome, we have:
$Y=\begin{cases}Y(1)&\text{if $\tau=1$ (treatment)}\\ Y(0)&\text{if $\tau=0$ (control)}.\end{cases}$
When $\tau=1$, $Y(0)$ is not observed: it is the counterfactual (Rubin, 1974).
In an RCT, assuming that $Y(1)$ and $Y(0)$ are independent from $\tau|X$, the conditional average treatment effect can be defined as:
$\mathbf{E}(Y(1)-Y(0)|X)=\mathbf{E}(Y|\tau=1,X)-\mathbf{E}(Y|\tau=0,X)$
with $X$ the characteristics of the patient.
More specifically, we compare $2$ FNNs that use specific loss functions to handle censored observations with IF, a specific type of random survival forest (RSF) that models interaction effects. These models are benchmarked with a penalized CoxPH with linear effects and the LASSO penalty.
3.2.1 Benchmark method: CoxPH with adaptive LASSO penalty
The Cox Proportional Hazards (CoxPH) model (Cox, 1972) is the most commonly used regression method in survival analysis. In a regression model, the objective is to estimate regression coefficients to assess the strength of association between the predictors $X$ and the outcome. For a given patient $i$, represented by a triplet $(X_{i},\tilde{T}_{i},D_{i})$, the hazard function $h(t,X_{i})$ examines how variables influence the rate of a particular event happening at a specific time $t$. It is written as:
$h(t|X_{i})=h_{0}(t)\exp(\beta^{T}X_{i}),\text{ for }i=1,\dots,n,$ (1)
where the baseline hazard function, $h_{0}(t)$, can be an arbitrary non-negative function of time. $\beta^{T}=(\beta_{1},\dots,\beta_{p})$ is the coefficient vector associated with the variables.
This CoxPH model relies on two assumptions. First, it supposes a linear relationship between the log hazard and the variables $X_{i}$, with the baseline hazard being an intercept term that varies with time. Additionally, the ratio of the instantaneous risk for any $2$ patients is independent of time $t$: this corresponds to the proportional hazard assumption.
With high-dimensional data, the model can become non-identifiable. A LASSO penalty can be introduced to select only the variables with the strongest effects on the outcome of interest. First introduced in the context of linear regression by Tibshirani (1996) and then adapted to survival analysis (Tibshirani, 1997), the LASSO penalty corresponds to the $L_{1}$ norm of the regression coefficients. The penalization term is denoted by:
[…]
The model’s degree of parsimony or complexity depends on the penalization parameter $\lambda$: it varies from 0 (complete model including all biomarkers) to $\infty$ (null model with no biomarkers). $\lambda$ is often estimated using cross-validation (CV).
In certain scenarios, LASSO can be inconsistent for variable selection. Here, we implement adaptive lasso (ALASSO) (Fan and Lv, 2008), a method based on an additive penalization term, which provides a more conservative choice of the shrinkage parameter $\lambda$. With ALASSO, adaptive weights are used for penalizing difference coefficients in the $L_{1}$ penalty. The penalization term becomes:
$pen(\lambda)=\lambda\left(\sum_{j=1}^{p}w_{j}|\beta_{j}|\right),$ (2)
where $w_{j}=\frac{1}{|\hat{\beta}_{j}|}$ are biomarker-specific weights, and $\hat{\beta}_{j},j=1,\dots,p$ are estimated by fitting a preliminary regular LASSO. The non-negative penalty parameter $\lambda$ for the adaptive lasso is chosen using the maximum cross-validated log-likelihood method.
3.2.2 Machine learning algorithms adapted to time-to-event outcomes
The hypotheses of the linear CoxPH model are restrictive, for instance, when nonlinear and complex interactions exist in the data. Here, we focus on artificial neural networks and random forests, as these algorithms are well-suited to model complex interaction patterns.
Kvamme et al. (2019) introduced a method called CoxCC that uses a loss function based on a case-control approximation. They proposed randomly sampling a new set of controls at each iteration instead of keeping the control samples fixed. The loss is written as:
$\mathcal{L}_{\text{CoxCC}}=\frac{1}{n}\sum\limits_{i:D_{i}=1}\log(\sum\limits_{j\in\tilde{\mathcal{R}}_{i}}\exp[\phi(X_{j})-\phi(X_{i})]),$ (3)
with $\tilde{\mathcal{R}}_{i}$ a subset of the risk set $\mathcal{\mathcal{R}}_{i}$ at time $t$ including individual $i$. $i$ represents the case, and the $j$’s are the controls sampled from the risk set. $\phi$ represents the transformation operated by the FNN on the input variables. CoxCC can be trained with this specific loss (Equation 3) using a mini-batch gradient descent algorithm.
The authors also developed a second version of their model, CoxTime, that is not constrained by the proportionality assumption. To do so, they added the time variable as an additional input to the model. The loss function can then be rewritten as:
$\mathcal{L}_{\text{Cox-Time}}=\frac{1}{n}\sum\limits_{i:D_{i}=1}\log\left(\sum\limits_{j\in\tilde{\mathcal{R}}^{i}}\exp\left[\phi(t_{i},X_{j})-\phi(t_{i},X_{i})\right]\right).$ (4)
In this paper, we compare the FNNs with IF (Hornung and Boulesteix, 2022), a type of random survival forest specifically designed to model quantitative and qualitative interaction effects in bivariate splits. $2$ variables are said to interact if the effect of one variable on the outcome depends on the value of the other variable. Here, we focus on biomarker-by-treatment interactions in the setting of an RCT. Two types of interactions are accounted for, quantitative and qualitative interactions, as categorized by Peto (1982). For a quantitative interaction, the treatment effect varies quantitatively depending on the biomarker value, but remains in the same direction, conditional on the biomarker value (e.g., the treatment effect increases as the biomarker value increases). Conversely, when considering a qualitative interaction, the direction of the treatment effect depends on the value of the biomarker.
3.2.3 Hyperparameter search
Specific hyperparameter search procedures are implemented for each machine learning model.
For the FNN models, we perform the hyperparameter search in the context of a 5-fold CV applied to the training set. Then the 5-fold CV is applied to the training set. It consists of randomly splitting the training data into $K=5$ folds. The model is trained using K1K-1 folds and validated on the remaining fold. The training is repeated until each fold is used as a validation set. Finally, the entire training set is used to train each network with the previously selected set of hyperparameters. We use the Tree-Parzen algorithm (Bergstra et al., 2011) to select hyperparameters iteratively in an informed manner. We define a search space for each hyperparameter with specific distribution and boundary values. These values are described in Table 6. A set of hyperparameters is randomly sampled, and the model is scored on each of the five validation folds. These five validation scores are averaged, and a new set of hyperparameters is sampled based on the value of the average score. The sampling of hyperparameter sets is repeated $200$ times.
Regarding the IF, it has been shown that the performance of random forests is relatively insensitive to changes in their hyperparameter values (Probst et al., 2019). Therefore, the default hyperparameter values are used in the implementation of IF. The number of variable pairs to sample for each split equals $\sqrt{p}/2$. The number of trees constituting each forest is set to $2,000$.
For the three case studies, we perform a double 5-fold CV on the entire dataset, because we have relatively small datasets and want to mimic an external test set. This strategy was first illustrated by Matsui et al. (2012). First, the real patient cohort is split into five folds, which constitute the outer loop. Then, we select one of the five folds as a test set and perform a 5-fold CV on the remaining data for each hyperparameter set: this is the inner loop. We choose the hyperparameter configuration with the minimum average validation loss obtained on the five folds of the inner loop. Finally, we fit the model with these optimal hyperparameters on the four folds of the outer loop and calculate predictions on the remaining test fold of the outer loop. This procedure is repeated on all the folds of the outer loop.
3.3 Data
The models are trained on two types of data: simulated data and real patient data sets. The synthetic data are generated to examine the operating characteristics and performances of the modeling techniques, while two high-dimensional patient clinical trial data sets are used for illustrative purposes.
3.3.1 Simulation study
In the first setting, we simulate nonlinear interactions between the biomarkers and the treatment variable using a full biomarker-by-treatment interaction model.
Rothwell (2005) put forward that the only reliable approach for assessing the predictiveness of biomarkers is to test their interaction with the treatment and introduce a full biomarker-by-treatment interaction model. Let $\tau$ denote the treatment indicator, with $\tau=1$ for patients receiving the treatment and $\tau=0$ for those in the control group, and let $X=(X_{1},\dots,X_{p})^{T}\in\mathbb{R}^{p}$ denote the vector of patient covariates. We can simulate data based on this model:
$h(t|\tau,X)=h_{0}(t)\exp\Bigg(\alpha_{\tau}\tau+\sum_{i=1}^{p}\beta_{i1}f_{i}(X_{i})+\sum_{i=1}^{p}\beta_{i2}f_{i}(X_{i})\tau\Bigg).$ (5)
In this model, we introduce nonlinearity through the choice of the function $f$, which enables the construction of complex interactions with the treatment. Our model extends the model from Haller et al. (2019), which considered the case of a single biomarker.
Here, $\alpha_{\tau}$ is chosen as $\alpha_{\tau}=ln(0.75)\approx-0.288$. We set $\beta_{i1}=0.288$, $f_{i}(X)=(2X-1)^{2}$ and $\beta_{i2}=-0.9$. The biomarkers are generated from a uniform distribution: $X_{i}\sim\mathcal{U}_{[0,1]}$. Since we intend to simulate an RCT, a total of $n$ patients per data set are randomly assigned with a 1:1 ratio to the experimental or control arm, with $P(\tau=1)=P(\tau=0)=\frac{1}{2}$.
Survival times are generated from a Weibull distribution. Thus, the risk can be written as follows:
$h(t|\tau,X)=b^{-a}at^{a-1}\exp\Bigg(\alpha_{\tau}\tau+\sum_{i=1}^{p}\beta_{i1}f_{i}(X_{i})\quad+\sum_{i=1}^{p}\beta_{i2}f_{i}(X_{i})\tau\Bigg)$ (6)
with $a$ the shape parameter and $b$ the scale parameter. For all scenarios, we generate exponential survival times (corresponding to a shape parameter $a=1$) with a time-constant baseline hazard rate. Using the inverse of the cumulative risk function for a Weibull distribution, we compute the theoretical survival probabilities (see details in Supplemental material):
$S(t|X)=\exp(-\int_{0}^{t}h(s|X)ds=\exp(-H_{0}(t)\exp(\beta X))$ (7)
A cohort of $n$ individuals ($n\in\{2,000;20,000\}$) is generated for each set, split into a $50\%$ training set and a $50\%$ test set. We generate independent censoring, and consider two censoring rates: a moderate censoring rate of $20\%$ and a high censoring rate of $50\%$. $q=20$ variables are generated for each set. Among the $q$ biomarkers, only $p=10$ are truly prognostic. For each scenario, $100$ replications are performed.
In the second data generation process, nonlinear interactions are simulated using Friedman’s random function generator (Friedman, 2001) within an accelerated failure time (AFT) model. Friedman’s random function generator enables us to generate random functions with second-order interactions between the survival data and the explanatory variables, as well as strong nonlinear effects. It is based on an AFT model that assumes a linear relationship between the logarithm of survival time $T$ and the variables. The random function generator is defined as:
$log\ T=m(X)+W\text{ with }W\sim\Gamma(2,1).$
A vector of variables $X=(X_{1},...,X_{20})$ is generated with $X\sim\mathcal{N}(0,I)$. $m(X)$ is defined by Friedman’s random function generator:
$m(X)=\sum_{l=1}^{10}a_{l}g_{l}(R_{l}).$
$\{a_{l}\}_{1}^{10}$ are randomly generated from a uniform distribution ($a_{l}\sim\mathcal{U}_{[-1,1]}$). $R_{l}$ is a random subset of the input vector $X$ of size $n_{l}=5$, implying high-order interaction effects. With Friedman’s function generator, the input variables are associated with the survival time at different levels:
$g_{l}(R_{l})=\exp\{-\frac{1}{2}(R_{l}-\mu_{l})^{T}V_{l}(R_{l}-\mu_{l})\}.$ (8)
Each mean vector $\{\mu_{l}\}_{1}^{10}$ is randomly generated with $\mu_{l}\sim\mathcal{N}(0,1)$. The matrix of variance-covariance $V_{l}$ is also randomly generated: $V_{l}=U_{l}D_{l}U_{l}^{T}$ with $U_{l}$ an orthonormal random matrix, $D_{l}=diag\{d_{l,1},...,d_{l,n_{l}}\}$ and $\sqrt{d_{lk}}\sim\mathcal{U}(0.1,2.0)$. Finally, we obtain the survival times by applying the exponential: $\exp(\log(T))$. This is the AFT-Friedman method.
Henderson et al. (2020) adapt the flexible random function generator specifically to treatment effects by introducing a second generator interacting with the treatment variable. For this purpose, the AFT-Friedman model can be rewritten as:
$\log T=m_{1}(X)+\tau\times m_{2}(X)+W\text{ with }W\sim\Gamma(2,1),$ (9)
where $m_{2}(X)$ is defined as :
$m_{2}(X)=\sum\limits_{i=1}^{10}a_{2l}g_{2l}(R_{2l}).$
The coefficients $a_{2l}$ are randomly generated from a uniform distribution $a_{2l}\sim\mathcal{U}_{[-0.2,0.3]}$. A total of $n$ patients per data set were randomly assigned $(1:1)$ to the experimental or control arm, with $P(\tau=1)=P(\tau=0)=\frac{1}{2}$.
By using the inverse of the cumulative risk function for a log-normal distribution, we are able to compute the theoretical survival probabilities (details in Supplemental material) :
$S(t|X)$ $=\exp\Bigg(-\int_{0}^{t}h(s|X)ds\Bigg)$
$=\exp\Bigg(-H_{0}(t\exp(m_{1}(X)+\tau\times m_{2}(X))\Bigg)$ (10)
A cohort of $n$ individuals ($n\in\{2,000;20,000\}$) is generated for each set, split into $50\%$ for the training set and $50\%$ for the test set. We generate independent censoring, and consider a moderate censoring rate of $20\%$. For each scenario, $100$ replications are performed.
3.4 Cases studies: application on cancer data sets
We apply the methods to gene expression data from a meta-analysis of neoadjuvant trials that included $614$ breast cancer patients treated by anthracyclines alone or anthracyclines plus taxanes (Ternès et al., 2018). We also use gene expression data from an RCT including $1,574$ patients, which evaluated the effect of adjuvant trastuzumab in early breast cancer (Pogue-Geile et al., 2013), and finally gene signature data an RCT in renal cancer with $n=886$ patients, which assessed first-line avelumab plus axitinib versus sunitinib in previously untreated patients with advanced renal cell carcinoma (Motzer et al., 2020).
3.4.1 Breast cancer study on taxane chemotherapy
We apply the models to a meta-analysis of neoadjuvant trials that included $614$ breast cancer patients treated by anthracycline-based chemotherapy with ($n=507$) or without ($n=107)$ taxane. The dataset comprises $p=1,689$ genes. This cohort is available in the biospear (Ternès et al., 2018) R package, which is publicly available.
Figure 1: Kaplan-Meier curves for the breast cancer cohort on taxane chemotherapy.
3.4.2 Breast cancer cohort with adjuvant trastuzumab
The models are also compared on data from an RCT funded by the National Cancer Institute and the National Institute of Health. This randomized phase III clinical trial evaluated the effect of adding trastuzumab to adjuvant chemotherapy on disease-free survival in early breast cancer patients with HER2+ tumors.
$n=1,574$ patients were included in the trial: $n=795$ patients received chemotherapy, while $n=779$ others received chemotherapy and trastuzumab. For each patient, the expression of $462$ genes is available, in addition to standard clinicopathological variables such as ER status or the number of positive nodes.
Figure 2: Kaplan-Meier curves for the breast cancer cohort with adjuvant trastuzumab.
3.4.3 Advanced Renal Cell Carcinoma cohort with avelumab
The third cohort analyzed comprises baseline tumor samples from the phase 3 JAVELIN Renal 101 trial, which assessed first-line avelumab plus axitinib versus sunitinib alone for progression-free survival in previously untreated patients with advanced renal cell carcinoma (Motzer et al., 2020). A total of $n=886$ patients were enrolled, with $n=443$ assigned to avelumab plus axitinib and $n=443$ to sunitinib. Consistent with the original study, we utilized the provided pathway scores, where each pathway score for a given sample was computed as the average of the standardized expression values of the genes comprising that pathway. The dataset features $58$ pathway scores alongside $2$ clinical variables: age and sex.
Figure 3: Kaplan–Meier survival curves for the advanced renal cell carcinoma cohort from the phase III JAVELIN Renal 101 trial.
3.5 Evaluation Measures
The models are compared using metrics adapted to the evaluation of treatment benefit: the Concordance index for benefit ($\text{C}_{\text{benefit}}$), the Integrated Calibration Index for benefit ($\text{ICI}_{\text{benefit}}$), the E50 for benefit ($\text{E50}_{\text{benefit}}$), and the root-mean-squared error (RMSE).
3.5.1 Observed and predicted treatment benefit
Since the counterfactual response for a given patient cannot be observed, the matched patient serves as a surrogate for this unobserved outcome. To construct such pairs, we use two matching procedures, based either on patients’ baseline covariates or on their predicted treatment benefit. For the first method, we compute a predicted benefit score for each individual. Then, within each arm (control and treatment), patients are ranked according to this score, and pairs are formed by aligning the two ranked lists so that patients with similar predicted benefit are matched. For the second method, we apply a covariate-based matching procedure based on patients’ characteristics and the Mahalanobis distance. For each treated patient $u$, we search among all control patients $v$ and identify the one whose covariate vector $X_{u}$ is closest to $X_{v}$ in terms of Mahalanobis distance. We first compare the two matching strategies on the simulation data, and subsequently retain and apply the predicted-benefit–based matching approach in the real-data analysis.
Importantly, we define the conditional average treatment benefit, $\hat{\theta}$, as an estimate of the CATE at time $t$, computed as the difference between the individual survival probability at a given time point with the treatment option being tested minus the probability with the control for the same patient. Given that matching is not perfect, the predicted benefits are averaged for the pair of patients previously matched, and the predicted benefit is defined as the average of the predicted benefits within each matched patient pair. We have :
$\hat{\theta}_{u,v}(t)$ $=\frac{[S_{u}(t|\tau=1,X_{u})-S_{u}(t|\tau=0,X_{u})]}{2}$ (11)
$+\frac{[S_{v}(t|\tau=1,X_{v})-S_{v}(t|\tau=0,,X_{v})]}{2},$
with $u$ the patient in the treated arm ($\tau=1$), and $v$ the patient in the control arm ($\tau=0$) for a given time $t$.
Observed treatment benefit $\tilde{\theta}$ is defined as the difference in time-to-event for $2$ matched patients. For each pair of patients ($u$ belonging to the control arm, and $v$ to the experimental arm), we calculate the observed treatment benefit using the following formula :
$\tilde{\theta}_{u,v}=\mathbf{1}_{Statut_{u}}\times\mathbf{1}_{\tilde{T}_{u}<\tilde{T}_{v}}-\mathbf{1}_{Statut_{v}}\times\mathbf{1}_{\tilde{T}_{v}<\tilde{T}_{u}}$ (12)
The first term of the equation $\mathbf{1}_{Statut_{i}}\times\mathbf{1}_{\tilde{T}_{u}<\tilde{T}_{v}}$ equals $1$ if the patient $i$ has a shorter time-to-event compared to patient $v$, and 0 otherwise. The reasoning is the same for the second term of the equation (for patient $v$). When the shortest survival time is censored, the treatment effect is not observed for the patient pair and equals 0.
Thus, for each pair of patients, the observed treatment benefit can take one of the following values: $1$ for observed treatment benefit, 0 for no observed treatment effect, and 1-1 for deleterious treatment effect.
3.5.2 Discrimination measures
The C-statistic at time $t$ measures the discrimination ability of a model, indicating its ability to distinguish high-risk and low-risk patients for a fixed time horizon. Specifically, it estimates the probability of agreement, i.e., the probability that $2$ patients randomly selected are ordered similarly in terms of survival prediction and in their observed survival data. Here, we use a concordance measure that accounts for the censored data using the inverse probability of censoring weighting (Uno et al., 2011). The concordance index for a horizon time $t$ is then:
$\hat{\text{C}}(t)=\frac{\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}D_{i}\hat{G}(\tilde{T}_{i})^{-2}\mathbb{I}\{\tilde{T}_{i}<\tilde{T}_{j},\tilde{T}_{i}<t\}\mathbb{I}\{\hat{S}(t|X_{i})<\hat{S}(t|X_{j})\}}{\sum\limits_{i=1}^{n}\sum\limits_{j=1}^{n}D_{i}\hat{G}(\tilde{T}_{i})^{-2}\mathbb{I}\{\tilde{T}_{i}<\tilde{T}_{j},\tilde{T}_{i}<t\}}$ (13)
The value of the C-statistic lies between 0 and $1$, with $0.5$ equivalent to a random prediction and $1$ corresponding to a perfect ability to rank. We compute the C-statistic for all individuals.
C-for-benefit ($\text{C}_{\text{benefit}}$) was introduced by Klaveren et al. (2017) as an extension of the C-statistic to measure treatment benefit at a horizon time $t$. It represents the probability that, among $2$ randomly chosen matched patient pairs, the pair with greater observed benefit also has a higher mean predicted benefit. It can be written as:
$\text{C}_{\text{benefit}}(t)=P(\hat{\theta}_{p_{k}}(t)>\hat{\theta}_{p_{l}}(t)|\tilde{\theta}_{p_{k}}(t)>\tilde{\theta}_{p_{l}}(t)),$ (14)
where $p_{k}$ and $p_{l}$ are $2$ pairs of patients.
3.5.3 Calibration measures
To evaluate the calibration capacity of the model, we use measures based on the calibration curve. Calibration commonly refers to the agreement between predicted and observed probabilities for the outcome (Austin et al., 2020). Here, we are interested in the correspondence between the predicted and observed treatment effects.
Calibration can be assessed by a smoothed calibration curve obtained through local regression of the observed pairwise treatment effect on the predicted pairwise treatment effect. The observed benefits are regressed on the predicted benefits using a locally weighted scatterplot smoother (loess). It is a non-parametric regression technique used to create a smoothing function by fitting a curve through a scatterplot of data points. It combines multiple local regressions to fit a curve to subsets of the data, allowing it to capture complex patterns without assuming a specific global model.
The Integrated Calibration-for-Benefit ($\text{ICI}_{\text{benefit}})$ (Rekkas et al., 2023; Maas et al., 2023) measures the average difference between predicted and smooth observed benefit across different levels of prediction determined by the empirical density function of the predicted pairwise treatment effect. $\text{E50}_{\text{benefit}}$ represents the $50^{th}$ percentile of the absolute distance between predicted and smooth observed benefit. It assesses how well the predicted treatment effect matches the observed treatment effect at the median. It offers insight into the model’s calibration at the central point rather than across the entire prediction range. For these measures, null values indicate perfect calibration, meaning that the model’s predicted treatment benefits align perfectly with the observed benefits.
The calibration quantile plot (Bouvier et al., 2024) represents the average observed benefit versus predicted benefit in quantiles of predicted benefit. The predicted benefit is divided into $5$ bins. In each bin, the average predicted benefit is compared to the observed benefit.
3.5.4 Root mean squared error
The predictive accuracy of the models is compared using the root mean squared error (RMSE) (Rekkas et al., 2023) adapted to time-to-event data. We compute RMSE for simulated data, where the true $\theta$ is known. We then have :
$\text{RMSE}=\sqrt{\frac{1}{n}\sum\limits_{i=1}^{n}(\theta_{i}(t)-\hat{\theta}_{i}(t))^{2}},$ (15)
with $\theta_{i}$ the true treatment benefit for individual $i$ at time $t$. For simulated data, $\hat{\theta}_{i}$ is the predicted benefit for individual $i$, that is, the difference in survival probabilities at a pre-specified time point $t$ for a patient with and without treatment given the covariate values:
$\hat{\theta}_{i}(t)=S_{i}(t|\tau=1,X_{i})-S_{i}(t|\tau=0,X_{i}).$
4 Results
4.1 Results on the simulation sets
4.1.1 Data generation process 1
The results reported here are the ones obtained using data simulated from the Weibull biomarker-by-treatment model with nonlinear interactions.
In Table 1 and Figure 5, the highest average $\text{C}_{\text{benefit}}$ were obtained with the IF model. It was also the closest to the oracle value. ALASSO obtained the lowest value in terms of $\text{C}_{\text{benefit}}$, which could indicate that the linear ALASSO model has difficulty capturing the treatment by biomarker interactions, whereas IF, due to its specific design for such interactions, demonstrates superior performance in modeling them.
$\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$ values remained substantially above zero across all models, which could highlight difficulties in obtaining solid calibration or limitations of the matching procedure between control group patients and treatment group patients. In contrast, RMSE values were not too far from zero, particularly for ALASSO. Note that for ALASSO, a small RMSE may not be so clinically useful if the C-for-benefit is close to 0.5.
If we only consider the machine learning methods, IF performed well in terms of discrimination performances and accuracy (higher concordance indices and lower RMSE), while FNNs had better calibration performances (lower $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$). Notably, CoxCC and CoxTime yielded similar results, as expected, since no interaction with time was generated in the simulation setup.
Table 1 also presents the results of models trained on an expanded cohort of patients ($n=10,000$). The results remain largely consistent between $n=1,000$ and $n=10,000$. For the machine learning methods, both the C-statistics and RMSE show only minor improvements compared to models trained on a smaller cohort. Meanwhile, $\text{C}_{\text{benefit}}$ decreases slightly, whereas $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$ exhibit slight increases.
Table 1: Average value at median time across the 100 simulation sets for data generation process 1.
$n=1,000$ $n=10,000$
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE
Oracle $\underset{(\pm 0.013)}{0.613}$ $\underset{(\pm 0.025)}{0.562}$ $\underset{(\pm 0.004)}{0.615}$ $\underset{(\pm 0.009)}{0.557}$
ALASSO $\underset{(\pm 0.014)}{0.567}$ $\underset{(\pm 0.028)}{0.504}$ $\underset{(\pm 0.062)}{0.373}$ $\underset{(\pm 0.059)}{0.373}$ $\mathbf{\underset{(\pm 0.026)}{0.118}}$ $\underset{(\pm 0.004)}{0.568}$ $\underset{(\pm 0.008)}{0.499}$ $\mathbf{\underset{(\pm 0.021)}{0.374}}$ $\mathbf{\underset{(\pm 0.021)}{0.376}}$ $\mathbf{\underset{(\pm 0.006)}{0.103}}$
CoxCC $\underset{\pm 0.018)}{0.545}$ $\underset{(\pm 0.043)}{0.546}$ $\underset{(\pm 0.117)}{0.321}$ $\underset{(\pm 0.116)}{\mathbf{0.321}}$ $\underset{(\pm 0.041)}{0.212}$ $\underset{(\pm 0.008)}{0.581}$ $\mathbf{\underset{(\pm 0.021)}{0.543}}$ $\underset{(\pm 0.0042)}{0.425}$ $\underset{(\pm 0.042)}{0.42}$ $\underset{(\pm 0.022)}{0.171}$
CoxTime $\underset{(\pm 0.016)}{0.543}$ $\underset{(\pm 0.04)}{0.541}$ $\underset{(\pm 0.107)}{\mathbf{0.320}}$ $\underset{(\pm 0.015)}{0.325}$ $\underset{(\pm 0.045)}{0.217}$ $\underset{(\pm 0.006)}{0.581}$ $\underset{(\pm 0.02)}{0.538}$ $\underset{(\pm 0.036)}{0.435}$ $\underset{(\pm 0.037)}{0.431}$ $\underset{(\pm 0.019)}{0.169}$
IF $\mathbf{\underset{(\pm 0.015)}{0.575}}$ $\underset{(\pm 0.04)}{\mathbf{0.561}}$ $\underset{(\pm 0.036)}{0.468}$ $\underset{(\pm 0.035)}{0.452}$ $\underset{(\pm 0.025)}{0.195}$ $\mathbf{\underset{(\pm 0.004)}{0.601}}$ $\underset{(\pm 0.011)}{0.539}$ $\underset{(\pm 0.015)}{0.473}$ $\underset{(\pm 0.013)}{0.479}$ $\underset{(\pm 0.008)}{0.164}$
  • The value in brackets is the standard deviation across the 100 simulation sets. The highest C values and lowest $\text{E}_{\text{benefit}}$, $\text{ICI}_{\text{benefit}}$ and RMSE values are in bold. Patients are matched on predicted benefit. $50\%$ censoring.
Figure 4: $\text{C}_{\text{benefit}}$ and RMSE values for each model in data generation process 1. Each boxplot displays $100$ points corresponding to each simulated set.
We also compared the models at different horizon points (Supplementary Material Table 7) and for different levels of censoring (Supplementary Material Table 8, Setting 1). All the results were improved when the censoring rate was lower ($20\%$ instead of $50\%$). Regarding the choice of the horizon time, the oracle values decreased over time, while the models performed similarly between the first and third quartiles of the time range.
The models were also compared using a matching procedure based on patients’ characteristics (Table 8, Setting 2). All the models performed worse compared to using a matching procedure based on predicted treatment benefit. As suggested by Klaveren et al. (2017), this is expected as the predicted treatment benefit for $2$ patients within a pair matched by patient characteristics tends to be less similar.
4.1.2 Data generation process 2
In Table 2, CoxCC, CoxTime, and IF exhibit comparable performances, with C values close to each other. CoxTime achieves the highest $\text{C}_{\text{benefit}}$ at $0.616$, with CoxCC and IF closely following. CoxTime also yields the lowest $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$, at $0.134$ and $0.153$. ALASSO consistently shows less good performance metrics compared to the other models.
Table 2 also presents the results for the model trained on a larger sample size. The results demonstrate only subtle differences across the evaluated metrics compared to the results obtained with the smaller sample size.
Table 2: Average value at median time across the 100 simulation sets for data generation process 2.
$n=1,000$ $n=10,000$
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE
Oracle $\underset{(\pm 0.038)}{0.841}$ $\underset{(\pm 0.047)}{0.583}$ $\underset{(\pm 0.039)}{0.838}$ $\underset{(\pm 0.039)}{0.584}$
ALASSO $\underset{(\pm 0.093)}{0.734}$ $\underset{(\pm 0.047)}{0.583}$ $\underset{(\pm 0.146)}{0.328}$ $\underset{(\pm 0.136)}{0.332}$ $\underset{(\pm 0.035)}{0.094}$ $\underset{(\pm 0.077)}{0.719}$ $\underset{(\pm 0.035)}{0.583}$ $\underset{(\pm 0.144)}{0.326}$ $\underset{(\pm 0.129)}{0.337}$ $\underset{(\pm 0.030)}{0.091}$
CoxCC $\mathbf{\underset{(\pm 0.048)}{0.876}}$ $\underset{(\pm 0.041)}{0.611}$ $\underset{(\pm 0.071)}{0.143}$ $\underset{(\pm 0.058)}{0.164}$ $\underset{(\pm 0.019)}{0.086}$ $\mathbf{\underset{(\pm 0.046)}{0.897}}$ $\mathbf{\underset{(\pm 0.039)}{0.599}}$ $\underset{(\pm 0.070)}{0.148}$ $\underset{(\pm 0.064)}{0.168}$ $\underset{(\pm 0.012)}{0.064}$
CoxTime $\underset{(\pm 0.057)}{0.868}$ $\mathbf{\underset{(\pm 0.042)}{0.616}}$ $\mathbf{\underset{(\pm 0.061)}{0.134}}$ $\mathbf{\underset{(\pm 0.053)}{0.153}}$ $\underset{(\pm 0.021)}{0.091}$ $\underset{(\pm 0.048)}{0.896}$ $\underset{(\pm 0.036)}{0.598}$ $\mathbf{\underset{(\pm 0.068)}{0.136}}$ $\mathbf{\underset{(\pm 0.060)}{0.153}}$ $\mathbf{\underset{(\pm 0.012)}{0.061}}$
IF $\underset{(\pm 0.051)}{0.875}$ $\underset{(\pm 0.037)}{0.608}$ $\underset{(\pm 0.090)}{0.146}$ $\underset{(\pm 0.079)}{0.172}$ $\mathbf{\underset{(\pm 0.021)}{0.081}}$ $\underset{(\pm 0.051)}{0.856}$ $\underset{(\pm 0.042)}{0.577}$ $\underset{(\pm 0.116)}{0.159}$ $\underset{(\pm 0.111)}{0.185}$ $\underset{(\pm 0.023)}{0.085}$
  • The value in brackets is the standard deviation across the 100 simulation sets. The highest C values and lowest $\text{E}_{\text{benefit}}$ and RMSE values are in bold. Patients are matched on predicted benefit. $50\%$ censoring.
Figure 5: $\text{C}_{\text{benefit}}$ and RMSE values for each model in data generation process 2. Each boxplot displays $100$ points corresponding to each simulated set.
4.2 Results on the three cancer data sets
FNN-based methods, IF and ALASSO were applied to $3$ cancer clinical trial data datasets. For the first two datasets, which involved early breast cancer, all comparison measures were computed on the test set at horizons of 2 and 5 years, as these are standard time points of interest for clinical investigators. The third dataset consisted of renal carcinoma patients, with comparison measures assessed at 1 year, as the endpoint is progression-free survival (PFS) in the advanced setting. We applied bootstrap resampling $M=100$ times to each dataset and computed confidence intervals for all evaluation measures.
4.2.1 Conditional average treatment effects of taxane chemotherapy
The bootstrapped evaluation measures at 2 and 5 years are presented in Table 3.
Overall, the neural network–based models adapted for survival, CoxCC and CoxTime, achieved the highest concordance values at 2-years, indicating superior discriminative ability compared to ALASSO and IF. When focusing on benefit-specific measures, IF slightly outperformed other models in terms of $\text{C}_{\text{benefit}}$, $\text{ICI}_{\text{benefit}}$, and $\text{E50}_{\text{benefit}}$, at both 2 and 5 years, suggesting it more accurately captures treatment effect heterogeneity. These results suggest that while neural networks are effective at predicting overall survival in this case study, they may not fully capture treatment interactions. In contrast, IF, designed to model individualized treatment effects, better identifies who benefits most from taxane chemotherapy.
Table 3: Bootstrapped evaluation measures at 2 and 5 years for the meta-analysis in the effect of taxane chemotherapy. - matching on predicted benefit.
$t=2$ $t=5$
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$
ALASSO $\underset{[0.499,0.563]}{0.511}$ $\underset{[0.514,0.750]}{0.650}$ $\underset{[0.038,0.212]}{0.127}$ $\underset{[0.063,0.206]}{0.138}$ $\underset{[0.15,0.905]}{0.563}$ $\underset{[0.491,0.758]}{0.632}$ $\underset{[0.032,0.19]}{\mathbf{0.099}}$ $\underset{[0.062,0.189]}{0.117}$
CoxCC $\underset{[0.632,0.714]}{\mathbf{0.673}}$ $\underset{[0.515,0.743]}{0.633}$ $\underset{[0.047,0.29]}{0.164}$ $\underset{[0.089,0.265]}{0.173}$ $\underset{[0.214,0.809]}{0.552}$ $\underset{[0.456,0.711]}{0.591}$ $\underset{[0.04,0.23]}{0.119}$ $\underset{[0.068,0.21]}{0.132}$
CoxTime $\underset{[0.629,0.713]}{\mathbf{0.673}}$ $\underset{[0.482,0.753]}{0.616}$ $\underset{[0.059,0.297]}{0.168}$ $\underset{[0.09,0.273]}{0.174}$ $\underset{[0.25,0.787]}{\mathbf{0.565}}$ $\underset{[0.460,0.692]}{0.577}$ $\underset{[0.045,0.23]}{0.123}$ $\underset{[0.062,0.218]}{0.131}$
IF $\underset{[0.454,0.577]}{0.506}$ $\underset{[0.554,0.760]}{\mathbf{0.656}}$ $\underset{[0.029,0.203]}{\mathbf{0.103}}$ $\underset{[0.047,0.19]}{\mathbf{0.109}}$ $\underset{[0.127,0.891]}{0.505}$ $\underset{[0.551,0.761]}{\mathbf{0.652}}$ $\underset{[0.033,0.174]}{\mathbf{0.099}}$ $\underset{[0.056,0.181]}{\mathbf{0.106}}$
  • The highest C values and lowest $\text{E}_{\text{benefit}}$, $\text{ICI}_{\text{benefit}}$ and RMSE values are in bold.
4.2.2 Conditional average treatment effects of adjuvant trastuzumab
For the second breast cancer data set, we observed in Table 4 that CoxCC and CoxTime reached the highest concordance values. Regarding treatment benefit, the machine-learning-based methods obtained higher values than ALASSO. CoxCC and CoxTime achieved the lowest $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$, indicating superior calibration of predicted benefit compared to ALASSO and IF. On the other hand, IF provided the highest $\text{C}_{\text{benefit}}$, suggesting it captures treatment heterogeneity more effectively, although at the cost of reduced calibration.
Table 4: Bootstrapped evaluation measures at 2 and 5 years in the RCT on adjuvant trastuzumab. - Matching on predicted benefit.
$t=2$ $t=5$
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$
ALASSO $\underset{[0.462,0.549]}{0.505}$ $\underset{[0.698,0.756]}{0.729}$ $\underset{[0.062,0.146]}{0.104}$ $\underset{[0.091,0.151]}{0.12}$ $\underset{[0.477,0.533]}{0.503}$ $\underset{[0.693,0.758]}{0.729}$ $\underset{[0.05,0.105]}{0.074}$ $\underset{[0.064,0.111]}{0.086}$
CoxCC $\underset{[0.55,0.625]}{\mathbf{0.588}}$ $\underset{[0.463,0.556]}{0.512}$ $\underset{[0.029,0.106]}{\mathbf{0.065}}$ $\underset{[0.036,0.097]}{0.069}$ $\underset{[0.546,0.607]}{\mathbf{0.581}}$ $\underset{[0.460,0.563]}{0.517}$ $\underset{[0.027,0.084]}{\mathbf{0.058}}$ $\underset{[0.039,0.086]}{0.063}$
CoxTime $\underset{[0.542,0.62]}{0.586}$ $\underset{[0.457,0.556]}{0.504}$ $\underset{[0.033,0.096]}{\mathbf{0.065}}$ $\underset{[0.037,0.094]}{\mathbf{0.064}}$ $\underset{[0.555,0.607]}{0.580}$ $\underset{[0.473,0.564]}{0.512}$ $\underset{[0.021,0.087]}{0.059}$ $\underset{[0.037,0.087]}{\mathbf{0.062}}$
IF $\underset{[0.468,0.535]}{0.503}$ $\underset{[0.715,0.774]}{\mathbf{0.744}}$ $\underset{[0.05,0.141]}{0.088}$ $\underset{[0.08,0.14]}{0.104}$ $\underset{[0.477,0.534]}{0.503}$ $\underset{[0.737,0.787]}{\mathbf{0.762}}$ $\underset{[0.04,0.091]}{0.069}$ $\underset{[0.052,0.099]}{0.076}$
  • The highest C values and lowest $\text{E}_{\text{benefit}}$, $\text{ICI}_{\text{benefit}}$ and RMSE values are in bold.
4.2.3 Conditional average treatment effects of avelumab
For the advanced renal cell carcinoma cohort (Table 5), CoxTime achieved the highest concordance at 1 year, indicating stronger discrimination overall, whereas IF obtained the highest $\text{C}_{\text{benefit}}$, suggesting a better ability to capture treatment heterogeneity. In terms of calibration of treatment benefit, CoxCC reached the lowest $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$, reflecting superior agreement between predicted and observed benefit. By contrast, IF showed higher calibration errors despite its strong discrimination.
Table 5: Bootstrapped evaluation measures at 1 year in the RCT on avelumab. - Matching on predicted benefit.
$t=1$
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$
ALASSO $\underset{[0.341,0.598]}{0.484}$ $\underset{[0.582,0.771]}{0.692}$ $\underset{[0.045,0.289]}{0.179}$ $\underset{[0.079,0.316]}{0.183}$
CoxCC $\underset{[0.417,0.651]}{0.544}$ $\underset{[0.51,0.61]}{0.561}$ $\underset{[0.022,0.14]}{\mathbf{0.061}}$ $\underset{[0.038,0.133]}{\mathbf{0.073}}$
CoxTime $\underset{[0.407,0.664]}{\mathbf{0.549}}$ $\underset{[0.514,0.612]}{0.563}$ $\underset{[0.021,0.144]}{0.068}$ $\underset{[0.036,0.125]}{0.078}$
IF $\underset{[0.39,0.639]}{0.517}$ $\underset{[0.695,0.759]}{\mathbf{0.728}}$ $\underset{[0.158,0.368]}{0.273}$ $\underset{[0.211,0.335]}{0.278}$
  • The highest C values and lowest $\text{E}_{\text{benefit}}$, $\text{ICI}_{\text{benefit}}$ and RMSE values are in bold.
5 Discussion
In this paper, we compared two FNN methods adapted to survival outcomes (CoxCC and CoxTime) and a random forest model (IF) for their ability to predict individualized treatment benefits in the context of time-to-event data, using a linear CoxPH model with an adaptive LASSO penalty as a benchmark method. The comparison was conducted on 2 simulation sets: one introducing nonlinear interaction with treatment and another based on Friedman’s random function generator with nonlinear and two-order interactions. Additionally, the models were evaluated on 3 case studies from oncology clinical trials.
The machine learning methods performed well in data generation process 1, especially in terms of discrimination capabilities for IF and calibration for FNNs. For data generation process 2, the performances of the machine learning methods were better than the linear ALASSO method with regard to the discrimination and calibration metrics, and RMSE.
Across the three clinical trial datasets, the models displayed complementary strengths in predicting treatment benefit. CoxCC and CoxTime generally achieved the lowest $\text{E50}_{\text{benefit}}$ and $\text{ICI}_{\text{benefit}}$, suggesting more reliable calibration of predicted benefit than IF. IF, on the other hand, tended to yield the highest $\text{C}_{\text{benefit}}$, indicating better identification of patients with larger differential treatment effects, though sometimes at the expense of calibration. ALASSO offered moderate, consistent performance across metrics. Overall, these findings reflect a trade-off: neural network–based models provide well-calibrated benefit estimates, while IF is more sensitive to heterogeneity, and ALASSO may offer a consistent alternative, even if it does not fit nonlinear effects or interactions in these types of cancer trial data sets.
In this paper, we used measures specifically adapted to treatment benefit evaluation in the context of time-to-event data. The assessment of treatment benefit prediction models has been an active area of research. Efthimiou et al. (2022) developed measures of discrimination and calibration, building upon the work of Klaveren et al. (2017), who introduced the C-for-benefit metric specifically for evaluating treatment benefits. We adapted their approach to time-to-event data by calculating this measure using survival probabilities. While C-for-benefit provides a flexible framework for comparing individualized treatment effects, its performance can vary with the matching procedure used and, as noted in recent methodological work (Efthimiou et al., 2025), it does not constitute a proper scoring rule. Additionally, we modified the RMSE metric to assess treatment benefit for time-to-event data by evaluating the difference between predicted and true treatment benefits, both derived from survival probabilities. Using adapted measures for treatment benefit is of crucial importance, as, for example, the standard concordance index is unable to directly capture the treatment benefit. Future work could include associating uncertainty measures with conditional average treatment effects, which could, for example, be evaluated by a bootstrap method as in Roblin et al. (2024).
Weberpals et al. (2025) emphasize the significant potential of leveraging complex causal machine learning models to refine individualized treatment effect estimation. Causal machine learning provides a promising framework for enhancing precision medicine by predicting patient-specific cancer treatment responses using rich, multimodal data. Although clinical applications of causal machine learning are still emerging, its capability to integrate high-dimensional data and provide robust causal inference shows great promise for improving personalized treatment choices. Our work is indeed centered on the RCT framework, which allows for the determination of causal relationships between an intervention and its outcome. However, the methods used in this paper could be adapted beyond RCTs and applied to real-world data cohorts by incorporating propensity scores or other causal approaches. Cui et al. (2023) proposed such an approach using causal forests, a machine learning method used to estimate heterogeneous treatment effects in observational studies.
Further work could focus on the relative contribution of the covariates, as it could enable us to identify specific biomarkers that predict larger treatment effects that could be further investigated in basic or clinical research project. Determining the relative importance of the different input variables is challenging in the case of FNNs. As it is a nonlinear type of model, simply examining the resulting weights and biases output by the model is not possible. Additionally, FNNs do not require the assumption of non-collinearity among the inputs. There could be informative interactions among groups of variables. To overcome this issue, different methodologies have been defined to assess variable contributions in FNNs that focus on the weights of the model (Garson, 1991) which could be explored. For the IF, we could evaluate the Effect Importance Measure. Feature importance could also be evaluated in both types of models using methods based on resampling and permutation Fisher et al. (2019).
6 Conclusion
Neural networks provide well-calibrated treatment benefit predictions, while random forests provide superior discrimination for identifying patients with heterogeneous responses. These complementary strengths reveal an inherent modeling trade-off. Our adapted metrics for survival benefits uncover performance differences that standard indices miss. These findings confirm causal machine learning’s promise for precision medicine, with methods applicable beyond clinical trials. Future work should quantify prediction uncertainty and identify key treatment-response biomarkers.
7 Acknowledgements
Elvire Roblin acknowledges financial support by Foundation Philanthropia Lombard-Odier. The authors also acknowledge the National Surgical Adjuvant Breast and Bowel Project (NSABP) investigators of the B-31 trial who submitted data from the original study to dbGaP (dbGaP Study Accession: phs000826.v1.p1) and the NIH data repository. The B-31 trial was supported by: National Cancer Institute, Department of Health and Human Services, Public Health Service, Grants U10-CA-12027, U10-CA-69651, U10-CA-37377, and U10-CA-69974, and by a grant from the Pennsylvania Department of Health.
8 Declaration of conflicting interests
The authors declare no potential conflict of interests.
Supplementary Material
A Hyperparameter search
Table 6: Hyperparameters search using Tree-Parzen Algorithm with the python package optuna.
Hyperparameter Value
Activation function $\{\text{elu, relu, tanh}\}$
Batch size $\{8,16,32,64,128,256\}$
Dropout rate $[0.01,0.5]$
$L_{2}$ regularization $[0.001,0.1]$
Learning rate $[0.001,0.01]$
Number of hidden layers $\{1,2,3,4\}$
Number of neurons per layer $[4,128]$
Optimization algorithm $\{\text{Adam, AdamAMSGRAD, RMSProp, SGDWR}\}$
B Simulated data
B.1 Details of computation
B.1.1 Data generation process 1
In this section, we introduce the calculation to obtain the theoretical survival values from the oracle model in the simulation data generation process 1.
To simulate survival times from a Weibull distribution, the following relationship can be used (Bender et al., 2005):
$T=H_{0}^{-1}[-\log(1-U)]\exp(\beta(X))$
with $U\sim\mathcal{U}[0,1]$. In the case of a Weibull risk function, $\lambda>0$ is the scale parameter, $\kappa$ is the shape parameter, and the cumulative risk function is written as:
$H_{0}(t)=\left(\frac{t}{\lambda}\right)^{\kappa}$
The inverse of $H_{0}$ is then written:
$H^{-1}_{0}(u)=\lambda u^{1/\kappa}$
We obtain:
$\boxed{T=\lambda[-\log(1-U)]^{1/\kappa}\exp(\beta X)}$
The basis risk function $h_{0}(t)$ is known and follows a certain probability distribution. In this case, we use the Weibull distribution. The risk function is then written as:
$h(t|X)=\frac{\kappa}{\lambda}\left(\frac{t}{\lambda}\right)^{\kappa-1}\exp(\beta X)$
with $h_{0}(t)=\frac{\kappa}{\lambda}(\frac{t}{\lambda})^{\kappa-1}$ the risk at baseline and $\beta$ the coefficients vector. We then obtain the survival function from the instantaneous risk :
$S(t|X)$ $=\exp(-\int_{0}^{t}h(s|X)ds$
$=\exp(-H_{0}(t)\exp(\beta X)).$
B.1.2 Data generation process 2
In this section, we introduce the calculation to obtain the theoretical survival values from the oracle model in the simulation data generation process 2.
To simulate survival times from an AFT model and a random generator based on Friedman’s function, the following relationship can be used (Leemis et al., 1990):
$T=\frac{H_{0}^{-1}[-\log(1-U)]}{\exp(m_{1}(X)+\tau\times m_{2}(X))},$
with $U\sim\mathcal{U}[0,1]$. In the case of a lognormal risk function, the cumulative risk function is written as :
$H_{0}(t)=-\log\left[1-\phi\left(\frac{\log(t)-\mu}{\sigma}\right)\right].$
The inverse of $H_{0}$ is then written:
$H_{0}^{-1}(u)=\exp(\sigma\phi^{-1}(1-exp(-u))+\mu).$
We obtain:
$\boxed{T=\frac{1}{\exp(m_{1}(X)+\tau\times m_{2}(X))}\exp(\sigma\phi^{-1}(U)+\mu)}.$
The basis risk function $h_{0}(t)$ is known and follows a certain probability distribution. In this case, we use the log-normal distribution. The risk function is then written as:
$h(t|X)=\exp(m_{1}(X)+\tau\times m_{2}(X))h_{0}(t\exp(m_{1}(X)+\tau\times m_{2}(X))),$
with $h_{0}(t)$ the risk at baseline and $\beta$ the coefficients vector. We then obtain the survival function from the instantaneous risk :
$S(t|X)$ $=\exp(-\int_{0}^{t}h(s|X)ds$
$=\exp(-H_{0}(t\exp(m_{1}(X)+\tau\times m_{2}(X))).$
B.2 Kaplan Meier Curves for data generation process 1 and 2
We can see the Kaplan-Meier curves of the treatment group and the control group for one of the dataset in data generation process 1 and 2 in the following figures.
Figure 6: Kaplan-Meier curve for one simulation set in data generation process 1, with $n=20000$.
Figure 7: Kaplan-Meier curve for one simulation set in data generation process 2, with $n=20000$.
B.3 Detailed results for data generation process 1
Table 7: Average value at the first and third quartiles of time across the 100 simulation sets for data generation process 1.
Q1 Q3
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE
Oracle $\underset{(\pm 0.021)}{0.619}$ $\underset{(\pm 0.022)}{0.604}$ $\underset{(\pm 0.013)}{0.613}$ $\underset{(\pm 0.024)}{0.504}$
ALASSO $\underset{(\pm 0.019)}{0.571}$ $\underset{(\pm 0.021)}{0.501}$ $\underset{(\pm 0.033)}{0.515}$ $\underset{(\pm 0.030)}{0.515}$ $\mathbf{\underset{(\pm 0.020)}{0.259}}$ $\underset{(\pm 0.013)}{0.569}$ $\underset{(\pm 0.020)}{0.503}$ $\underset{(\pm 0.044)}{0.440}$ $\underset{(\pm 0.042)}{0.439}$ $\mathbf{\underset{(\pm 0.013)}{0.095}}$
CoxCC $\underset{(\pm 0.017)}{0.543}$ $\underset{(\pm 0.045)}{0.536}$ $\underset{(\pm 0.137)}{0.395}$ $\underset{(\pm 0.134)}{0.394}$ $\underset{(\pm 0.029)}{0.309}$ $\underset{(\pm 0.017)}{0.543}$ $\mathbf{\underset{(\pm 0.039)}{0.538}}$ $\underset{(\pm 0.114)}{0.338}$ $\underset{(\pm 0.112)}{0.342}$ $\underset{(\pm 0.026)}{0.118}$
CoxTime $\underset{(\pm 0.016)}{0.543}$ $\underset{(\pm 0.039)}{0.537}$ $\mathbf{\underset{(\pm 0.113)}{0.388}}$ $\mathbf{\underset{(\pm 0.127)}{0.388}}$ $\underset{(\pm 0.033)}{0.312}$ $\underset{(\pm 0.016)}{0.543}$ $\underset{(\pm 0.046)}{0.534}$ $\mathbf{\underset{(\pm 0.111)}{0.336}}$ $\mathbf{\underset{(\pm 0.110)}{0.339}}$ $\underset{(\pm 0.029)}{0.124}$
IF $\mathbf{\underset{(\pm 0.016)}{0.571}}$ $\mathbf{\underset{(\pm 0.043)}{0.590}}$ $\underset{(\pm 0.051)}{0.499}$ $\underset{(\pm 0.057)}{0.471}$ $\underset{(\pm 0.018)}{0.307}$ $\mathbf{\underset{(\pm 0.013)}{0.575}}$ $\underset{(\pm 0.040)}{0.541}$ $\underset{(\pm 0.036)}{0.479}$ $\underset{(\pm 0.031)}{0.466}$ $\underset{(\pm 0.014)}{0.099}$
  • The value in brackets is the standard deviation across the 100 simulation sets. The highest C values and lowest E for benefit and RMSE values are highlighted in bold. Patients are matched on predicted benefit. There is $50\%$ censoring.
Table 8: Average value at median time across the 100 simulation sets for data generation process 1.
Setting 1 Setting 2
Model C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE C $\text{C}_{\text{benefit}}$ $\text{E50}_{\text{benefit}}$ $\text{ICI}_{\text{benefit}}$ RMSE
Oracle $\underset{(\pm 0.013)}{0.615}$ $\underset{(\pm 0.028)}{0.619}$ $\underset{(\pm 0.028)}{0.559}$
ALASSO $\underset{(\pm 0.013)}{0.569}$ $\underset{(\pm 0.032)}{0.501}$ $\underset{(\pm 0.035)}{0.684}$ $\underset{(\pm 0.034)}{0.682}$ $\mathbf{\underset{(\pm 0.020)}{0.144}}$ $\underset{(\pm 0.02)}{0.503}$ $\underset{(\pm 0.042)}{0.442}$ $\underset{(\pm 0.04)}{0.443}$ $\mathbf{\underset{(\pm 0.025)}{0.141}}$
CoxCC $\underset{(\pm 0.017)}{0.550}$ $\underset{(\pm 0.046)}{0.568}$ $\mathbf{\underset{(\pm 0.155)}{0.475}}$ $\mathbf{\underset{(\pm 0.148)}{0.477}}$ $\underset{(\pm 0.037)}{0.212}$ $\underset{(\pm 0.036)}{0.528}$ $\mathbf{\underset{(\pm 0.0099)}{0.350}}$ $\mathbf{\underset{(\pm 0.01)}{0.349}}$ $\underset{(\pm 0.041)}{0.214}$
CoxTime $\underset{(\pm 0.019)}{0.551}$ $\mathbf{\underset{(\pm 0.045)}{0.575}}$ $\underset{(\pm 0.151)}{0.483}$ $\underset{(\pm 0.143)}{0.484}$ $\underset{(\pm 0.038)}{0.206}$ $\underset{(\pm 0.034)}{0.531}$ $\underset{(\pm 0.084)}{0.360}$ $\underset{(\pm 0.084)}{0.357}$ $\underset{(\pm 0.046)}{0.219}$
IF $\mathbf{\underset{(\pm 0.013)}{0.590}}$ $\underset{(\pm 0.053)}{0.562}$ $\underset{(\pm 0.027)}{0.694}$ $\underset{(\pm 0.027)}{0.679}$ $\underset{(\pm 0.023)}{0.173}$ $\mathbf{\underset{(\pm 0.036)}{0.539}}$ $\underset{(\pm 0.032)}{0.479}$ $\underset{(\pm 0.037)}{0.488}$ $\underset{(\pm 0.025)}{0.195}$
  • The value in brackets is the standard deviation across the 100 simulation sets. The highest C values and lowest E for benefit and RMSE values are highlighted in bold. Patients are matched on predicted benefit.
  • Setting 1: Matching on predicted benefit. There is $20\%$ censoring.
  • Setting 2: Matching on covariates.There is $50\%$ censoring.
Figure 8: Calibration plot for one dataset in data generation process 1 and with CoxTime predictions. The left figure displays all the points, and the right figure shows quintiles.